OpenStudy (anonymous):

HELP !! Giving Out Medals ;) y = 3/7x + 11 –3x + 7y = 13 A.No, since the y-intercepts are different. B.Yes, since the slopes are the same and the y-intercepts are the same. C.No, since the slopes are different. D.Yes, since the slopes are the same and the y-intercepts are different.

OpenStudy (studygurl14):

What's the question? Like, what are you trying to find out...?

OpenStudy (studygurl14):

whether or not they are parallel...?

OpenStudy (anonymous):

yes! @StudyGurl14

OpenStudy (studygurl14):

-3x + 7y = 13 Move the x value over 7y = 3x + 13 divide to isolate y y = 3/7 x + 13/7 So..what's the answer?